# Online Makespan Scheduling under Scenarios

### v1

@@ 3 Near-Tight Bounds for $m=2$ Machines
A significant portion of the analysis is focused on proving that if line 9 is executed, Invariant 3.5 is indeed maintained. We refer the reader to the full version of the paper for further details.
@@
Proof 3.7.
To ease the notation once again, let $X\coloneqq 3+\sqrt{17}$.
@@
Notice that the aforementioned family of instances fails to achieve small competitive ratios for reasons similar to those that require Property (iii) in Invariant 3.5.
It can also be shown through a simple array of instances that, no algorithm that satisfies Rule 3.3 can beat the bound of $\nicefrac{{5}}{{3}}$, for which we refer the reader to the full version of the paper.
4 Tight Bounds for Sufficiently Many Scenarios
One of our main results (Theorem 1.2) states that for any number $m\in\mathbb{N}$ of machines, there exists a number $K$ of scenarios such that no deterministic algorithm for OMSS$(m,K)$ can have a competitive ratio less than the trivial upper bound $m$, even for unit processing times.
Due to space constraints, we merely sketch a proof for the special case of $m=3$. While doing so, we showcase key ideas of the general argument, although we forfeit the restriction onto hypertrees for the sake of a more clearly structured description. In the context of scheduling, this is indeed not a significant change.
@@ 4 Tight Bounds for Sufficiently Many Scenarios
For OMHC$(3)$, there exists no $r$-competitive online algorithm for any $r<3$, even when restricted to instances with at most $233$ hyperedges of size at most $3$ each.
@@
The problem OMHC$(m)$, which we formulated in Section 1.2, is equivalent to the unit processing time case of OMSS$(m,K)$ with $K$ part of the input, so that the lower bounds are immediately implied for the latter as well.
@@
To prove Theorem 4.1, we provide a family of instances on which any online algorithm is forced to output a monochromatic hyperedge of size $3$. For each of the instances, we also maintain (offline) colorings $c\colon[j]\to\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$, with the property that if two nodes $v,v^{\prime}$ are contained in a common hyperedge, we have $c(v)\neq c(v^{\prime})$. The latter coloring suggests an offline optimal value of $1$ for each of the instances, implying the competitive ratio of $3$ immediately. The main difference between the online and offline colorings is, as expected, that we may change the offline coloring completely at all times as long as its makespan equals $1$. Among other modifications, we often employ a permutation $\pi\colon\{\mathrm{red,blue,yellow}\}\to\{\mathrm{red,blue,yellow}\}$ on connected components to realize this possibility. We must trace the offline coloring simultaneously because the hyperedges are often defined adversarially according to both colorings of the current hypergraph.
@@
To differentiate between the online and offline colorings, we denote by $J_{i}\subseteq V$ the subset of nodes assigned to the online color $i$ for $i\in\{1,2,3\}$. Right before the revelation of a node $j$, we implicitly mean $J_{i}\cap[j-1]$ whenever we mention $J_{i}$ for some $i\in\{1,2,3\}$. Inspired by our original scheduling problem, we call the objective value $\max_{e\in E}\max_{i\in\{1,2,3\}}|e\cap J_{i}|$ of an instance its makespan. Without loss of generality, we truncate an instance as soon as the online solution is forced to obtain a makespan of $3$.
Our gadget: Seven-node subhypergraph $\mathcal{S}$.
@@
Initially, we study a subhypergraph $\mathcal{S}$ that we use as a building block throughout our construction. The hypergraph $\mathcal{S}$ is given by the nodes $V(\mathcal{S})=\{1,\ldots,7\}$ and hyperedges
@@
The hyperedges have size at most $3$ and the set $E(\mathcal{S})$ is closed under taking subsets of size at least $2$, i.e., if $e\in E(\mathcal{S})$, $|e|=3$ and $e^{\prime}\subseteq e$ with $|e^{\prime}|=2$, then $e^{\prime}\in E(\mathcal{S})$. The hypergraph $\mathcal{S}$ together with the maximal hyperedges $e_{1},\ldots,e_{5}$ is given in Figure 3.
$e_{1}$$e_{2}$$e_{3}$$e_{4}$$e_{5}$3333111122225555444466667777
We first assume that the seven nodes in $V(\mathcal{S})$ are assigned to exactly two out of the three online colors. Since the online colors are initially symmetric, let us assume that we use the online colors $1$ and $2$. By a careful case distinction, we may observe:
@@
For any partitioning $V(\mathcal{S})=J_{1}\dot{\cup}J_{2}$ of the nodes that admits a makespan of at most $2$, there exists a node coloring $c\colon V(\mathcal{S})\to\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$ such that there are two hyperedges $e\in E(\mathcal{S}[J_{1}]),e^{\prime}\in E(\mathcal{S}[J_{2}])$ with $c(v)\in\{\mathrm{red},\mathrm{blue}\}$ (or another subset of $\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$ of size $2$) for all $v\in e\cup e^{\prime}$.
In other words, we have induced monochromatic hyperedges of size $2$ with two different colors, both with respect to the online coloring. Applying this repeatedly, we can further force the algorithm to a monochromatic hyperedge of size $3$, which we show next.
Pigeonholing the two-color case.

### v2

@@ 3.2 The Algorithm
3.3 Analysis of Algorithm 1
It is evident that if Line 7 is executed, the assumption is maintained and we obtain a $\nicefrac{{5}}{{3}}$-competitive schedule. It is also easy to observe that the execution of Line 3 provides this: In the first scenario, the second machine is a least loaded machine, so that $\rho$ or $\alpha$ cannot increase. [^2] This is straightforward to observe and left to the reader as an exercise. Moreover, if the schedule was previously dominated by the first machine, the third property is maintained as we do not increase $p(J_{1}\cap S_{1})$.
Finally, we deal with the execution of Line 13: the proxy competitive ratio does not exceed $\nicefrac{{5}}{{3}}$ after the last single-scenario job, which we prove below.
Lemma 3.13.
For an instance of OMSS(2,2) satisfying the properties in Observation 3.9, let $j\in[n]$ be the largest index with $j\in S_{1}\triangle S_{2}$. Then, if Algorithm 1 satisfies Invariant 3.12 on the first $j$ jobs, then it also satisfies Invariant 3.12 on the first $\ell$ jobs for $\ell\in\{j+1,\ldots,n-1\}$.
Proof 3.14.
  1. (i)
    $\alpha\leq 2$: We claim that $\alpha$ stays constant after the insertion of the $j$-th job. Indeed, $p(J_{2}\cap S_{2}\cap[\ell])>p(J_{2}\cap S_{1}\cap[\ell])$ is maintained (as $p(J_{2}\cap S_{2}\cap[\ell])-p(J_{2}\cap S_{1}\cap[\ell])$ is constant). Similarly, both $p(J_{1}\cap S_{2}\cap[\ell])$ and $p(J_{2}\cap S_{1}\cap[\ell])+p(J_{1}\cap S_{1}\cap[\ell])-p(J_{2}\cap S_{2}\cap[\ell])$ are constant in $\ell\in\{j+1,\ldots,n\}$.
  2. (ii)
    $\rho\leq\nicefrac{{5}}{{3}}$: By Lemma 3.10, it suffices to observe that the anticipation prior to the assignment of the jobs $\ell\in\{j+1,\ldots,n\}$ is constant.
  3. (iii)
    Third property: If there was no machine dominance after the insertion of the $j$-th job, machine dominance does not emerge since the jobs $\ell\in\{j+1,\ldots,n-1\}$ are all assigned to the second machine. Otherwise, the condition of the third property is maintained for the same reason.
This lemma does not cover the $n$-th job, however, the $n$-th job only needs to satisfy $\rho\leq\nicefrac{{5}}{{3}}$, which follows by the fact that $\alpha\leq 2$ prior to its assignment as we have seen, and Lemma 3.10.
All that remains to analyze is that if Line 9 is executed, Invariant 3.12 is maintained. Notice that machine dominance cannot occur as the prior makespan is still among the top two largest loads and is in the opposite machine of the makespan. Hence the third property does not need to be checked.
We apply a case distinction. In each case, we either bound the proxy competitive ratio $\rho\leq\nicefrac{{5}}{{3}}$ and the anticipation $\alpha\leq 2$ by means of inequalities, or find inequalities that are necessary for a condition to fail and show that conjunction of these conditions has no solutions. As all inequalities are linear in $p(J_{i}\cap S_{k}\cap[j-1])$ and $p_{j}$, the reader might interpret the remainder of the proof as a sequence of cheerful linear inequalities which they are encouraged to trace. Accordingly, we ease the notation, and define:
$x_{1}$ $\coloneqq p(J_{1}\cap S_{1}\cap[j-1])$
$x_{2}$ $\coloneqq p(J_{1}\cap S_{2}\cap[j-1])$
$x_{3}$ $\coloneqq p(J_{2}\cap S_{1}\cap[j-1])$
$x_{4}$ $\coloneqq p(J_{2}\cap S_{2}\cap[j-1])$
$x_{5}$ $\coloneqq p_{j}$
We start with the case that $j$ has a large processing time.
Case 1: $x_{5}\geq 2x_{1}-x_{2}$.
In this case, the attempt to assign the job $j$ to the first machine failed because
$p(J_{1}\cap S_{2}\cap[j])=x_{2}+x_{5}\geq x_{2}+2x_{1}-x_{2}=2x_{1},$
so that the assignment to the second machine yields a makespan of
$p(J_{2}\cap S_{2}\cap[j])=x_{4}+x_{5}\geq x_{4}+2x_{1}-x_{2}\geq x_{1}.$
The proxy competitive ratio is at most
$\rho\leq\max\left\{\frac{x_{4}+x_{5}}{x_{5}},\frac{2x_{4}+2x_{5}}{x_{2}+x_{4}+x_{5}}\right\}\leq\frac{5}{3}.$
Indeed, if $x_{1}\geq 2x_{2}$, then we have
$\frac{x_{4}+x_{5}}{x_{5}}\leq\frac{x_{4}+2x_{1}-x_{2}}{2x_{1}-x_{2}}\leq\frac{3x_{1}-x_{2}}{2x_{1}-x_{2}}\leq\frac{2.5x_{1}}{1.5x_{1}}=\frac{5}{3},$
which is equivalent to $2x_{5}\geq 3x_{4}$. Otherwise we have
$\frac{2x_{4}+2x_{5}}{x_{2}+x_{4}+x_{5}}\leq\frac{5x_{4}}{x_{2}+2.5x_{4}}\leq\frac{5x_{1}}{0.5x_{1}+2.5x_{1}}\leq\frac{5}{3}.$
Finally, the anticipation is bounded by
$\alpha\leq\frac{x_{4}+x_{5}}{x_{2}+x_{4}+x_{5}-x_{1}}\leq\frac{x_{5}}{x_{2}+x_{5}-x_{1}}\leq\frac{2x_{1}-x_{2}}{x_{2}+2x_{1}-x_{2}-x_{1}}=\frac{2x_{1}-x_{2}}{x_{1}}\leq 2,$
so that all properties in Invariant 3.12 are shown.
In the sequel, we assume that $x_{5}\leq 2x_{1}-x_{2}$ and distinguish some cases based on how $x_{1},x_{2},x_{3},x_{4}$ relate to each other.
Case 2: $x_{1}\geq x_{2}\geq x_{3}$ and $x_{2}\geq x_{4}$.
The $j$-th job was assigned to the second machine. This cannot be because of properties (ii) or (iii), as these could only be violated if $x_{5}\geq 2x_{1}-x_{2}$, which we have already excluded. Hence, assigning the $j$-th job to the first machine must have created a proxy competitive ratio of larger than $\frac{5}{3}$, i.e.,
$\frac{2x_{2}+2x_{5}}{x_{2}+x_{4}+x_{5}}>\frac{5}{3}\Leftrightarrow x_{2}+x_{5}>5x_{4}.$ (12)
If the makespan is still attained by $J_{1}\cap S_{1}$, then the proxy competitive ratio $\rho$ has not increased and is bounded by $\frac{5}{3}$ by induction. Otherwise the new makespan is $x_{4}+x_{5}\geq x_{1}$ and the proxy ratio is bounded by $\rho\leq\frac{x_{4}+x_{5}}{\max\{x_{5},\frac{x_{2}+x_{4}+x_{5}}{2}\}}\leq\frac{3}{2}$, similar to the proof of Graham’s List Scheduling Algorithm since $j$ was assigned to a least loaded machine.
We only need to check the property (ii), i.e., that $\alpha\leq 2$. However, the anticipation is bounded by $\frac{x_{1}}{x_{2}}\leq 2$ since the initial schedule must satisfy the property (iii).
Case 3: $x_{1}\geq x_{3}$ and $x_{4}\geq x_{2}$.
We claim that this case never occurs, that is, in this case the job $j$ is always assigned to the first machine. Since we assume $x_{5}\leq 2x_{1}-x_{2}$, the property (iii) cannot be the reason the assignment to the first machine has failed. Similarly, if $x_{5}+x_{2}\geq x_{1}$, the anticipation is bounded by $\frac{x_{5}+x_{2}}{x_{1}}\leq\frac{2x_{1}-x_{2}+x_{2}}{x_{1}}=2$, and otherwise the anticipation cannot increase, either; so the property (ii) is also maintained while assigning to the first machine. Finally, we note that by $x_{4}\geq x_{2}$, the first machine is least loaded in the second scenario, therefore by the proof of Graham’s List Scheduling, the proxy ratio either decreases or is bounded by $\frac{3}{2}$.
Case 4: $x_{1}\geq x_{3}\geq x_{2}\geq x_{4}$, proof by LP solver.
Exactly as above, the assignment cannot have failed because of the properties (ii) and (iii). Therefore, assigning $j$ to the first machine must yield a proxy competitive ratio larger than $\frac{5}{3}$, which implies both $2x_{5}\leq 3x_{2}$ and $x_{5}+x_{2}\leq 5x_{4}$.
As $x_{4}\leq x_{2}$, the second machine is least loaded in the second scenario and therefore $\rho\leq\frac{5}{3}$ is secured if we place $j$ to the second machine. By the above observation, we can further bound the anticipation by
$\alpha\leq\min\left\{\frac{x_{4}+x_{5}}{x_{3}+(x_{4}+x_{5}-x_{1})},\frac{x_{1}}{x_{2}+(x_{1}-x_{4}-x_{5})}\right\}.$
If $x_{4}+x_{5}\geq x_{1}$, the first term is valid, which means
$\frac{x_{4}+x_{5}}{x_{3}+(x_{4}+x_{5}-x_{1})}>2\Leftrightarrow 2(x_{1}-x_{3})>x_{4}+x_{5}.$ (13)
Gathering all constraints, we obtain system of inequalities
$x_{1}$ $\geq x_{3}$ (14)
$x_{3}$ $\geq x_{2}$ (15)
$x_{2}$ $\geq x_{4}$ (16)
$2x_{5}$ $\leq 3x_{2}$ (17)
$x_{4}+x_{5}$ $\geq x_{1}$ (18)
$2x_{1}-2x_{2}$ $\geq x_{5}+x_{4}$ (19)
$x_{4}+x_{5}$ $\geq 2x_{3}$ (20)
$x_{2}+x_{5}$ $\geq 5x_{4}$ (21)
The only feasible solutions to this system of inequalities are of the type $(4x_{4},2x_{4},2x_{4},x_{4},3x_{4})$. For these solutions, it is immediate that $\alpha\leq 2$, from which the claim follows.
Similarly if $x_{4}+x_{5}\geq x_{1}$, the inequality (18) is reversed and the inequality (19) is replaced by $\frac{x_{1}}{x_{3}+(x_{1}-x_{4}-x_{5})}>2\Leftrightarrow 2x_{4}+2x_{5}\geq 2x_{3}+x_{1}$, so that all schedules to be inspected are the solutions of the following system:
$x_{1}$ $\geq x_{3}$ (22)
$x_{3}$ $\geq x_{2}$ (23)
$x_{2}$ $\geq x_{4}$ (24)
$2x_{5}$ $\leq 3x_{2}$ (25)
$x_{4}+x_{5}$ $\leq x_{1}$ (26)
$2x_{4}+2x_{5}$ $\geq 2x_{3}+x_{1}$ (27)
$x_{4}+x_{5}$ $\geq 2x_{3}$ (28)
$x_{2}+x_{5}$ $\geq 5x_{4}$ (29)
This system of inequalities has the same set of solutions as well, so that the anticipation is bounded by $\alpha\leq 2$.
As all cases are covered, we conclude that Algorithm 1 maintains the properties in Invariant 3.12 and is therefore $\nicefrac{{5}}{{3}}$-competitive.
3.4 Lower Bounds
@@
Proof 3.16.
To ease the notation once again, let $X\coloneqq 3+\sqrt{17}$. We demonstrate our family of instances by means of a case distinction, that is, our instances mostly agree with each other in earlier jobs and their later jobs are revealed depending on the schedule.
@@
One might wonder how the gap between the upper bound in Theorem 1.1 and the lower bound in Theorem 3.15 can be narrowed down. We can observe through a simple family of instances that, if that is possible at all by means of improving the upper bound, then dropping Rule 3.3 is necessary.
Observation 3.17.
Let $r<\nicefrac{{5}}{{3}}$ be arbitrary. There exists no deterministic $r$-competitive algorithm for OMSS(2,2) which satisfies Rule 3.3.
Proof 3.18.
Consider an instance of $n=5$ jobs with processing times $p_{1}=p_{2}=p_{4}=1$, $p_{3}=2$, $p_{5}=3$. The scenarios are given by $S_{1}=\{1,2,4,5\}$, $S_{2}=\{3,4,5\}$.
The first two jobs must be assigned to opposite machines; for otherwise one could truncate the instance, i.e., consider the restricted instance of the first $n=2$ jobs with $p_{j}\equiv 1$, $S_{1}=\{1,2\}$, $S_{2}=\emptyset$ instead. On this restricted instance, the algorithm would be no better than $2$-competitive. Hence, we may assume without loss of generality that $1\in J_{1}$ and $2\in J_{2}$.
The third job is assigned to the first machine without loss of generality. Then, as $4\in S_{1}\cap S_{2}$ and the first machine uniquely admits makespan, we must assign the fourth job to the second machine, and similarly, the fifth job is assigned to the second machine as well.
In total, the makespan is given by
$p(J_{1}\cap S_{2})=1+1+3=5,$
although an optimal solution is given by $J_{1}=\{1,2,3,4\}$ and $J_{2}=\{5\}$, which has a makespan of $3$.
[…]
Figure 5: The online schedule in the proof of Observation 3.17 and the offline optimum.
In the above setting, the right course of action would have indeed been to violate Rule 3.3 and assign $4\in J_{1}$. Hence, it is justified to suspect that the algorithm can be improved by altering this rule according to the sizes of jobs $j\in S_{1}\cap S_{2}$, although as seen in Theorem 3.15, it can only improve the ratio by less than $0.027$.
4 Tight Bounds for $m=3$ and Sufficiently Many Scenarios
One of our main results (Theorem 1.2) states that for any number $m\in\mathbb{N}$ of machines, there exists a number $K$ of scenarios such that no deterministic algorithm for OMSS(m,K)(m,K) can have a competitive ratio less than $m$, even for unit processing times.
@@ 4 Tight Bounds for $m=3$ and Sufficiently Many Scenarios
Let $K\geq 233$. Then for Online Makespan Hypergraph Coloring(m)(m), there exists no $r$-competitive online algorithm for any $r<3$, even when the problem is restricted to $3$-uniform instances with at most $K$ hyperedges.
@@
The problem Online Makespan Hypergraph Coloring(m)(m), which we formulated in Section 1.2, is equivalent to the unit processing time case of OMSS(m,K)(m,K) with $K$ part of the input, so that the lower bounds are immediately implied for the latter as well for $K\geq 233$.
@@
To prove Theorem 4.1, we provide a family of instances on which any online algorithm is forced to output a monochromatic hyperedge of size $3$. For each of the instances, we also maintain (offline) colorings $c\colon[j]\to\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$ to be feasible, i.e., with the property that if two nodes $v,v^{\prime}$ are included in a common hyperedge $e$, we have $c(v)\neq c(v^{\prime})$. The latter coloring suggests an offline optimum of value $1$ for each of the instances, implying the competitive ratio of $3$ immediately. The main difference between the online and offline colorings is, as expected, that we may change the offline coloring completely at all times as long as it remains feasible. Among other modifications, we often employ a permutation $\pi\colon\{\mathrm{red,blue,yellow}\}\to\{\mathrm{red,blue,yellow}\}$ to realize this possibility. We must trace the offline coloring simultaneously because the hyperedges are often defined adversarially according to both colorings of the current hypergraph.
@@
In the following, let $J_{i}\subseteq V$ denote the subset of nodes assigned to the online color $i$ for $i\in\{1,2,3\}$. Right before the insertion of a node $j$, we implicitly mean $J_{i}\cap[j-1]$ whenever we mention $J_{i}$ for some $i\in\{1,2,3\}$. Inspired by our original scheduling problem, we call the objective value $\max_{e\in E}\max_{i\in\{1,2,3\}}|e\cap J_{i}|$ of an instance its makespan. Without loss of generality, we can truncate an instance as soon as it obtains a makespan of $3$.
4.1 Our gadget: Seven-node subgraph $\mathcal{S}$
@@
Initially, we study a subgraph $\mathcal{S}$ that we use as a building block throughout our construction. The hypergraph $\mathcal{S}$ is given by the nodes $V(\mathcal{S})=\{1,\ldots,7\}$ and hyperedges
$E(\mathcal{S})=\big\{$ $\{1,2\},\{1,3\},\{2,3\},\{1,2,3\},\{2,4\},\{4,5\},\{2,5\},\{2,4,5\},$
$\{4,6\},\{6,7\},\{4,7\},\{4,6,7\},\{3,6\},\{3,7\}\big\}.$
One can verify that the hyperedges have size at most $3$ and the hyperedge set $E(\mathcal{S})$ is closed under taking subsets of size at least $2$, i.e., if $e\in E(\mathcal{S})$, $|e|=3$ and $e^{\prime}\subseteq e$ with $|e^{\prime}|=2$, then $e^{\prime}\in E(\mathcal{S})$. The hypergraph $\mathcal{S}$ together with the maximal hyperedges $e_{1},\ldots,e_{5}$ is given in Figure 6.
$e_{1}$$e_{2}$$e_{3}$$e_{4}$$e_{5}$33112255446677
We first assume that the seven nodes are assigned to exactly two out of three online colors. Since the online colors are initially symmetric, let us assume that we use the online colors $1$ and $2$.
Assumption 4.2.
$e_{k}\cap J_{3}=\emptyset$ for all $k\in\{1,2,3\}$.
Moreover, for brevity, we only mention the revelation of inclusion-wise maximal hyperedges whenever we reveal a copy of this gadget, although the subsets of these hyperedges of size $2$ are being introduced as well.
First, we describe a case distinction about the introduction of the gadget, always satisfying Assumption 4.2. We reveal the first three nodes that appear in the common hyperedge $e_{1}=\{1,2,3\}$. If all three nodes were assigned to a common online color, we would obtain a makespan of $3$, which ultimately proves the theorem. Therefore, we assume that the nodes were partitioned in two online colors; as the nodes are symmetric so far, we may assume without loss of generality that the first two nodes were assigned to the first color and the third job to the second, see Figure 7.
@@
Figure 7: First three nodes.
Now we introduce two more nodes $4,5$ in this order one by one and reveal that
$e_{2}=\{2,4,5\}.$
If both nodes $4$ and $5$ are assigned to the second online color, we may assign the remaining two nodes $6$ and $7$ arbtirarily and ignore them from now on, see Figure 8.
@@
Figure 8: The five relevant nodes if $\tau(4)=\tau(5)=2$.
Otherwise, one of the nodes $4$ and $5$, without loss of generality, the node $4$, is assigned to the first color. We reveal the sixth node along with the hyperedges
$e_{3}\supseteq\{4,6\},e_{4}\supseteq\{3,6\}.$
If the sixth node is assigned to the second color, we may ignore the seventh node similar to above, see Figure 9.
[…]
Figure 9: The six relevant nodes if $\tau(6)=2$.
If the sixth node is assigned to the first online color, we reveal the seventh node along with the hyperedges
$e_{3}=\{4,6,7\},e_{5}\supseteq\{3,7\}.$
Throughout the remainder of this section, we let $\mathcal{S}$ denote the resulting hypergraph with seven nodes and fourteen hyperedges (including subsets).
Since both the sixth and the seventh nodes are assigned to the first online color, the seventh node is forced to be assigned to the second online color. Our final assignment looks as in Figure 10.
[…]
Figure 10: The hypergraph $\mathcal{S}$ together with the online and offline colorings, under the assumption that $\tau(6)=1$.
@@
For any partitioning $V(\mathcal{S})=J_{1}\dot{\cup}J_{2}$ of the nodes that admits a makespan of at most $2$, there exists a node coloring $c\colon V(\mathcal{S})\to\{\operatorname{\mathrm{blue},\mathrm{red},\mathrm{yellow}}\}$ such that there are two hyperedges $e\in E(\mathcal{S}[J_{1}]),e^{\prime}\in E(\mathcal{S}[J_{2}])$ with $c(e)=c(e^{\prime})=\{\mathrm{red},\mathrm{blue}\}$.
Proof 4.3.
See Figures 8, 9 and 10 for the offline coloring in the respective cases. The hyperedges $(e,e^{\prime})$ in question are $(\{1,2\},\{4,5\})$; $(\{2,4\},\{3,6\})$ and $(\{2,4\},\{3,7\})$, respectively.
In other words, we have induced monochromatic hyperedges of both colors with respect to the online coloring. Applying this repeatedly, we can force the algorithm to a monochromatic hyperedge of size $3$, which we show next.
4.2 Pigeonholing the two-color case